# Projected and Solvable Topological Heavy Fermion Model of Twisted Bilayer Graphene

### v1

@@ VII.1 Appendix: Mapping an interacting Green’s function to a non-interacting Green’s function
$\rho_{ij}^{\Sigma}(\omega)=-\frac{1}{2\pi i}\left[\Sigma_{ij}(\omega+i0^{+})-\Sigma_{ji}^{*}(\omega+i0^{+})\right]$ (89)
@@
$\rho_{ij}^{\Sigma}(\omega)\approx\sum_{n}\bigg{(}v_{ij}^{n}\bigg{)}^{2}\delta(\omega-\epsilon_{ij}^{n})$ (91)
where $v_{ij}^{n},\epsilon_{n}$ are introduced to describe the behaviors of $\rho_{ij}^{\Sigma}$. From Eq. 89, we also note that
$v_{ij}^{n}=(v_{ji}^{n})^{*},\quad\quad\epsilon_{ij}^{n}=\epsilon_{ji}^{n}$ (92)
This also indicates that the self-energy can be written as
$\Sigma_{ij}(i\omega_{n})=\sum_{n}\frac{(v_{ij}^{n})^{2}}{i\omega_{n}-\epsilon_{ij}^{n}}$ (93)
@@
$\hat{H}^{aux}=\sum_{ij}t_{ij}\gamma_{i}^{\dagger}\gamma_{j}+\sum_{nij}\epsilon^{n}_{ij}a_{nij}^{\dagger}a_{nij}+\sum_{nij}v_{ij}^{n}\bigg{[}a_{nij}^{\dagger}\gamma_{i}+\gamma_{j}^{\dagger}a_{nij}\bigg{]}$ (94)
We can now calculate the corresponding Green’s functions with respect to the Hamiltonian $\hat{H}^{aux}$. The Green’s function (which is a matrix) is then
$G^{aux}(i\omega_{n})=\langle T_{\tau}\begin{bmatrix}\gamma(\tau)&a(\tau)\end{bmatrix}\cdot\begin{bmatrix}\gamma^{\dagger}(0)\\ a^{\dagger}(0)\end{bmatrix}\rangle$ (95)
where $\gamma(\tau)=[\gamma_{1}(\tau),...]$ is the vector formed by all the $\gamma$ electron operators, and $a(\tau)=[a_{111}^{\dagger}(\tau),...]$ is the vector formed by all the $a$ electron operators. The Green’s function then can be calculated as
$G^{aux}(i\omega_{n})=\rangle\begin{bmatrix}i\omega_{n}\mathbb{I}-t&-{V}\\ -V^{\dagger}&i\omega_{n}\mathbb{I}-\epsilon\end{bmatrix}^{-1}$ (96)
@@
$V_{i^{\prime},nji}=v_{ji}^{n}\delta_{i^{\prime},i},$ (97)
$t$ denotes the hopping matrix with element $t_{ij}$ for row $i$ and column $j$, and $\epsilon$ is a diagonal matrix with diagonal elements $\epsilon^{n}_{ij}$. We now aim to get the Green’s function of $\gamma$. We first note the following fact
@@
$G^{aux,\gamma}_{ij}(i\omega_{n})=\langle T_{\tau}\gamma_{i}(\tau)\gamma_{j}(0)\rangle=\begin{bmatrix}i\omega_{n}\mathbb{I}-t-\tilde{\Sigma}(i\omega_{n})\end{bmatrix}^{-1}_{ij}$ (99)
@@
$\tilde{\Sigma}_{ij}(i\omega_{n})=\bigg{[}V\cdot[i\omega_{n}\mathbb{I}-\epsilon]^{-1}\cdot V\bigg{]}_{ij}=\sum_{n}\frac{v_{ij}^{n}v_{ij}^{n}}{i\omega_{n}-\epsilon_{ij}^{n}}$ (100)
@@
$\tilde{\Sigma}_{ij}(i\omega_{n})=\Sigma_{ij}(i\omega_{n})$ (101)
@@
VIII Appendix: Berry curvature and band gap in a two-orbital toy model
@@
$u[\mathbf{k}_{0}+(p_{x},p_{y})]=\frac{1}{\sqrt{2\sqrt{X^{2}+Y^{2}+Z^{2}}(\sqrt{X^{2}+Y^{2}+Z^{2}}-Z})}\begin{bmatrix}Z-\sqrt{X^{2}+Y^{2}+Z^{2}}&X+iY\end{bmatrix},\quad$
@@
Here, we could observe that $\Omega(\mathbf{k}_{0})$ is proportional to $1/|E_{\mathbf{k}_{0},+}|^{2}$. To obtain a diverged Berry curvature, we need to either let $|v_{x}[\mathbf{k}_{0}]v_{y}[\mathbf{k}_{0}|\rightarrow\infty$, or let $|E_{\mathbf{k}_{0},+}|^{2}\rightarrow 0$. However, we notice that (from Sections VIII, 110 and 105),
@@
$\approx$ ±$\pm\sqrt{(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{x}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{x}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{y}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{y}[\mathbf{k}_{0}]p_{y})^{2}}$
$\approx$ ±$\pm(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})+O(p_{x}^{2},p_{y}^{2},p_{x}p_{y})$ (118)

### v2

@@ VIII.3 Mapping an interacting Green’s function to a non-interacting Green’s function
$\rho_{ij}^{\Sigma}(\omega)=-\frac{1}{2\pi i}\left[\delta\Sigma_{ij}(\omega+i0^{+})-\delta\Sigma_{ji}^{*}(\omega+i0^{+})\right]$ (106)
@@
$\rho_{ij}^{\Sigma}(\omega)\approx\sum_{n}v_{i}^{n}(v_{j}^{n})^{*}\delta(\omega-\epsilon_{n})$ (108)
where $v_{i}^{n},\epsilon_{n}$ are introduced to describe the behaviors of $\rho_{ij}^{\Sigma}$. This also indicates that the self-energy can be written as
$\delta\Sigma_{ij}(i\omega)=\sum_{n}\frac{v_{i}^{n}v_{j}^{n,*}}{i\omega-\epsilon_{n}}$ (109)
@@
$\hat{H}^{aux}=\sum_{ij}\tilde{t}_{ij}\gamma_{i}^{\dagger}\gamma_{j}+\sum_{n}\epsilon^{n}a_{n}^{\dagger}a_{n}+\sum_{n}\bigg[v_{i}^{n}\gamma_{i}^{\dagger}a_{n}+\text{h.c.}\bigg]$ (110)
where $\tilde{t}=t-N$ which contains both the non-interacting contributions and the static (non-decay) contributions from the self-energy. We can now calculate the corresponding Green’s functions with respect to the Hamiltonian $\hat{H}^{aux}$. The Green’s function (which is a matrix) is then
$G^{aux}(\tau)=\langle T_{\tau}\begin{bmatrix}\gamma(\tau)&a(\tau)\end{bmatrix}\cdot\begin{bmatrix}\gamma^{\dagger}(0)\\ a^{\dagger}(0)\end{bmatrix}\rangle$ (111)
where $\gamma(\tau)=[\gamma_{1}(\tau),$...] is the vector formed by all the $\gamma$ electron operators, and $a(\tau)=[a_{1}^{\dagger}(\tau),$...] is the vector formed by all the $a$ electron operators. The Green’s function then can be calculated as
$G^{aux}(i\omega_{n})=\begin{bmatrix}i\omega_{n}\mathbb{I}-\tilde{t}&-{V}\\ -V^{\dagger}&i\omega_{n}\mathbb{I}-\epsilon\end{bmatrix}^{-1}$ (112)
@@
$V_{i,n}=v_{i}^{n}$ (113)
and $\epsilon$ is a diagonal matrix with diagonal elements $\epsilon_{n}$. We now aim to get the Green’s function of $\gamma$. We first note the following fact
@@
$G^{aux,\gamma}_{ij}(i\omega_{n})=\langle T_{\tau}\gamma_{i}(\tau)\gamma^{\dagger}_{j}(0)\rangle=\begin{bmatrix}i\omega_{n}\mathbb{I}-\tilde{t}-\tilde{\Sigma}(i\omega_{n})\end{bmatrix}^{-1}_{ij}$ (115)
@@
$\tilde{\Sigma}_{ij}(i\omega_{n})=\bigg[V\cdot[i\omega_{n}\mathbb{I}-\epsilon]^{-1}\cdot V\bigg]_{ij}=\sum_{n}\frac{v_{i}^{n}(v_{j}^{n})^{*}}{i\omega_{n}-\epsilon_{n}}$ (116)
@@
$\tilde{\Sigma}_{ij}(i\omega_{n})=\delta\Sigma_{ij}(i\omega_{n})$ (117)
@@
IX Appendix: Berry curvature and band gap in a two-orbital toy model
@@
$u[\mathbf{k}_{0}+(p_{x},p_{y})]=\frac{1}{\sqrt{2\sqrt{X^{2}+Y^{2}+Z^{2}}(\sqrt{X^{2}+Y^{2}+Z^{2}}-Z})}\begin{bmatrix}Z-\sqrt{X^{2}+Y^{2}+Z^{2}}&X+iY\end{bmatrix}$$\quad$
@@
Here, we could observe that $\Omega(\mathbf{k}_{0})$ is proportional to $1/|E_{\mathbf{k}_{0},+}|^{2}$. To obtain a diverged Berry curvature, we need to either let $|v_{x}[\mathbf{k}_{0}]v_{y}[\mathbf{k}_{0}|\rightarrow\infty$, or let $|E_{\mathbf{k}_{0},+}|^{2}\rightarrow 0$. However, we notice that (from Sections IX, 126 and 121),
@@
$\approx$ $\pm\sqrt{(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{x}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{x}[\mathbf{k}_{0}]p_{y})^{2}+(v_{x}[\mathbf{k}_{0}]{n}^{y}_{x}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{y}[\mathbf{k}_{0}]p_{y})^{2}}$
$\approx$ $\pm(E_{\mathbf{k}_{0},+}+v_{x}[\mathbf{k}_{0}]{n}^{x}_{z}[\mathbf{k}_{0}]p_{x}+v_{y}[\mathbf{k}_{0}]n_{y}^{z}[\mathbf{k}_{0}]p_{y})+O(p_{x}^{2},p_{y}^{2},p_{x}p_{y})$ (134)